Find two consecutive integers such that half the smaller number plus three times the larger number is -32

CAN YOU PLEASE USE THE GUESS METHOD IF POSSIBLE
THANK YOU

Answer by Boreal(15235)

I don't use the specific GUESS method but here is my approach.
both these integers are less than 0.
let x be the smaller integer and x+1 the larger one. That means x+1 is closer to 0.
half the smaller plus 3 times the larger becomes (1/2)x+3(x+1)=-32
(1/2)x+3x+3=-32
3.5x=-35
divide by 3.5
x=-10
and x+1=-9
half the smaller is -5 (remember that -10 is a smaller number) added to 3 times the larger (-9=-27) equals -32.